# NotebookOS: A Notebook Operating System for Interactive Training with On-Demand GPUs

### v1

@@ Abstract.
Interactive notebook programming is universal in modern ML (machine learning) and AI (artificial intelligence) workflows. Notebook software like Jupyter and Google Colab provides a user-friendly, interactive, web-based programming interface and is widely used across science and engineering domains. A dominant application of production notebook workloads is interactive deep learning training (IDLT). To guarantee high interactivity, modern notebook platforms typically reserve GPU resources within actively running notebook sessions. These notebook sessions are long-running but exhibit intermittent and sporadic GPU usage. Consequently, during most of their lifetimes, notebook sessions do not use the reserved GPUs, resulting in extremely low GPU utilization and prohibitively high cost.
In this paper, we introduce NotebookOS, a GPU-efficient notebook platform designed to meet the unique requirements of IDLT. NotebookOS uses a replicated notebook kernel design, where each kernel consists of three replicas distributed across separate GPU servers and synchronized via Raft. To optimize GPU utilization, NotebookOS oversubscribes server resources via kernel replication to leverage the relatively high task inter-arrival times in IDLT workloads. By dynamically allocating GPUs to kernel replicas only while they are actively executing notebook cells, NotebookOS maximizes the likelihood of immediate and interactive training upon notebook notebook-cell task submission. NotebookOS also migrates kernel replicas and automatically scales the GPU cluster under overload conditions. We evaluate NotebookOS extensively using production notebook workloads. Evaluation results show that NotebookOS saves 1,187+ GPU hours over a 17.5-hour real-world IDLT workload while greatly enhancing interactivity.
@@ 1. Introduction
Traditional GPU-based, batch deep learning training (BDLT) features batch-style, long-running workloads (tiresias_nsdi19, ; gandiva_osdi18, ; optimus_eurosys18, ; themis_nsdi20, ; philly_trace_atc19, ; alibaba_cluster_2020_trace_nsdi22, ) that require uninterrupted access to GPU resources over long periods of time, such as large-scale model training. These traditional systems focus on maximizing throughput and job-completion time (JCT), often prioritizing long-running tasks that can afford latency and resource contention. In contrast, notebook-oriented, interactive deep learning training (IDLT) workloads—which consist of tasks like model and program debugging, dynamic/iterative model adjustments, hyperparameter tuning—demand a different set of performance characteristics. This set includes low latency, responsiveness, high interactivity, and efficient management of many short-running tasks. The short-lived tasks that dominate IDLT workloads are executed within the context of long-lived user sessions, such as Jupyter Notebook sessions.
@@
To guarantee interactivity, today’s notebook services typically provision and reserve GPU resources during the entire lifetime of a Notebook session. These notebook sessions are long-running, but with fragmented cell executions, IDLT workloads exhibit intermittent, sporadic, and often transient GPU usage. Users simply spend more time developing a notebook than they do executing the notebook’s tasks. As a result, notebook sessions spend a majority of their time not using the reserved GPUs. Such dynamic usage patterns lead to extremely low GPU utilization and prohibitively high cost.
@@
We build NotebookOS, a novel and first-of-its-kind notebook platform designed for the unique requirements of IDLT. NotebookOS is distinct in both purpose and design from traditional GPU cluster computing systems (gandiva_osdi18, ; themis_nsdi20, ; tiresias_nsdi19, ; gavel_osdi20, ; antman_osdi20, ; pollux_osdi21, ; sia_sosp23, ). Specifically, NotebookOS prioritizes interactivity, low response time, and efficient management of numerous short-lived GPU tasks. By addressing the unique needs of users actively working in notebooks—where intermittent, short-lived tasks are the norm—NotebookOS realizes a novel notebook platform that tailors fine-grained, flexible GPU allocation to effectively meet the demands of short and dynamic GPU tasks.
@@
At NotebookOS’s core is a collection of novel techniques and design tenets that enable the efficient support of interactive machine learning training workloads. NotebookOS uses a pluggable notebook task scheduling and placement mechanism whose default policy is custom-designed to maximize interactivity. Specifically, NotebookOS uses a replicated kernel design in which each kernel is composed of three replicas placed on separate hosts. By dynamically allocating GPUs to kernels only while they are actively executing code, NotebookOS maximizes the likelihood that a kernel will be able to begin training immediately upon code submission. Likewise, NotebookOS also oversubscribes host resources to take advantage of the fact that IDLT workloads have a relatively high task inter-arrival time (IAT). This maximizes kernel availability and minimizes wait times during user code submission.
@@
    Comprehensive evaluation of NotebookOS using prototype and simulation on production notebook workloads.
We aim to pave the way for more interactive and user-friendly GPU-enabled environments for ML/AI practitioners.
@@ 2. Background and Motivation
In this section, we first present an overview of relevant background material beginning with a discussion of Jupyter Notebook and Notebook-as-a-Service (NaaS). We then present a workload analysis that motivates the need for a system tailored for notebook IDLT workloads.
Figure 1. Client-server Jupyter Notebook architecture.
@@ 2.1. Jupyter Notebook
Jupyter Notebook (jupyter_notebook, ) is an open-source, web-based interactive development environment (IDE) designed for interactive programming. Jupyter Notebook is commonly used with Python, though there is also support for other programming languages, including R (jupyter_irkernel, ) and Julia (jupyter_ijulia, ), among others (jupyter_available_kernels, ). Jupyter Notebook has become the de facto development environment for data-driven programming: ML/AI, data science, and data analytics. This is due in large part to its versatility, ease-of-use, and interactivity.
Terminology. Notebooks are interactive documents that combine code, text for explaining the code, and visualizations for displaying the code’s inputs and outputs. These elements are contained within the cells of the notebook, which are the basic “units” of a notebook. When a user runs code in some cell(s), the code is sent to a separate process, called a kernel, which resides either on the user’s local computer or on a remote server. The kernel executes the code and returns results back to the notebook document where they are displayed in the cell(s). We refer to this process as a cell task execution. A notebook session is a persistent, working instance of a Jupyter Notebook environment where the state of variables, imports, and other execution context are maintained and reused by the associated kernel. Finally, we define an IDLT task as a cell task execution of GPU operations—such as model training and inference—during an interactive workload that involves debugging and testing, initial model design, and other similar activities. See Figure 1 for an illustration.
@@ 2.2. Interactive Deep Learning Training
(a) Training task duration CDF.
(b) Inter-arrival time (IAT) CDF.
(c) GPU util. CDF (Sensei).
(d) GPU and CPU usage timeline (Sensei).
Figure 2. Workload characteristics of three representative GPU cluster traces.
@@ 2.3. Real-World IDLT Workload Analysis
Real-world IDLT workloads exhibit distinct behaviors compared to traditional long-running training jobs. In order to better understand these workloads and their requirements, we present a workload analysis of the production notebook IDLT traces collected from Adobe Sensei (adobe_sensei, ), an industry-leading AI service platform. Sensei hosts containerized notebook sessions within AWS EC2 GPU virtual machine (VM) instances. Sensei also manages the GPU resource allocations for the notebook sessions. To deploy a notebook container, the user specifies how many GPUs should be allocated. Once the notebook container is launched, GPUs are bound to the notebook session and the user can interact with the back-end notebook kernel (i.e., an IPython process) via a web-based Jupyter Notebook interface.
The IDLT traces (SenseiTrace) include 545,467 individual training events collected over a ten-month period from January–October, 2021. The workload spans 2,826 p3.16xlarge GPU VMs, 3,648 containers, and over 166 million GPU hours. We focus on a representative subset spanning June 1–August 31. The sample granularity of this trace is 15 seconds.
To provide context, we compare SenseiTrace with two other widely used, publicly available GPU cluster traces. The first trace is the Alibaba GPU Cluster’20 trace (alibaba_cluster_2020_trace_nsdi22, ) (AlibabaTrace). This trace captures a mix of training and inference jobs running state-of-the-art ML algorithms. Data were collected from a large production cluster with 6,500 GPUs spread across 1,800 servers, between July and August 2020.
The second trace in our comparison is the Philly GPU trace (philly_trace_atc19, ) (PhillyTrace). This trace is a 6.6GB representative subset of first-party deep learning training (DLT) workloads from Microsoft’s internal Philly clusters. It includes a total of 117,325 individual jobs collected from a cluster of 2,490 GPUs distributed across 552 servers between August 7, 2017, and December 22, 2017.
We chose these two traces for comparison because they represent typical DLT workloads, which are characterized by long-running training tasks that are scheduled by a batch scheduler and span several hours. As we show in our analysis, the workload captured by SenseiTrace differs considerably from that of PhillyTrace and AlibabaTrace.
@@ 2.3.1. Training Task Duration
Figure 2(a) displays a cumulative distribution function (CDF) of the training task durations, in seconds, of DLT tasks submitted by users to the Sensei GPU cluster. The data used to plot the CDF were taken from the June, July, and August 2021 trace dataset. 50% of user-submitted tasks are 2 minutes or less, while 75% are 5 minutes or less. 90% of training tasks are 17 minutes or less, 95% are 36 minutes or less, and the duration of the 99th percentile training task is 182 minutes (i.e., $\bar{3}$ hours). Based on this, it is clear that an overwhelming majority of training tasks are short-lived.
The SenseiTrace workload consists of a significantly high fraction of short training tasks compared to both PhillyTrace and AlibabaTrace. From this graph, we can see that the 50th percentile of task durations is 120 seconds (2 minutes), 621 seconds (10.35 minutes), and 957 seconds (15.95 minutes) for SenseiTrace, PhillyTrace, and AlibabaTrace, respectively. This observation highlights the difference between traditional BDLT workloads (i.e., PhillyTrace and AlibabaTrace) and notebook IDLT workloads (i.e., SenseiTrace). Note that a meaningful comparison of task durations less than 15 seconds is impossible due to the 15-second granularity of SenseiTrace.
@@
2.3.2. Training Task Inter-Arrival Time
Figure 2(b) plots a CDF of the IATs of training tasks from the Sensei GPU cluster traces. For all three traces, the IATs were measured within each user session independently, rather than the cluster-wide IATs of tasks submitted by any active user session, to ensure a fair comparison. From this graph, we can see that the 50th percentile of task IATs is 300 seconds (5 minutes), 44 seconds, and 38 seconds for SenseiTrace, PhillyTrace, and AlibabaTrace, respectively.
The observed SenseiTrace IATs align with the nature of IDLT workloads, where users intermittently make small, debugging-like changes to their models or Python code before submitting a short cell task to test the changes. In particular, SenseiTrace exhibits significantly longer IATs for more than 50% of tasks compared to PhillyTrace and AlibabaTrace. This is likely due to SenseiTrace users engaging in more extensive interactive cycles, iteratively testing and refining their code before each submission, rather than executing fully tested, production-ready DLT tasks scheduled by a batch GPU scheduler.
@@ 2.4. Insight and Challenges
This behavior leads to extremely low GPU utilization, as discussed in section 2.3.
@@
$\bullet$ C3: Resource Elasticity. Low resource utilization stems from static GPU over-provisioning and a lack of elastic resource management in NaaS platforms. Sensei, operating roughly 12k V100 GPUs across 3k p3.16xlarge EC2 instances, incurs roughly \$18.3 million in monthly costs even with long-term reservation discounts. This high TCO (total cost of ownership), combined with the extremely low resource utilization, highlights the urgent need for a GPU-efficient notebook platform that can flexibly adapt based on real-time workload requirements.
@@ 3. NotebookOS Design
Workload Requirements. We present NotebookOS, a GPU-efficient notebook platform to address the challenges outlined in section 2. Traditional GPU cluster schedulers are designed with long-running GPU-intensive tasks in mind. They often prioritize job completion time (JCT) (gandiva_osdi18, ; optimus_eurosys18, ; tiresias_nsdi19, ), throughput (gavel_osdi20, ), and fairness (themis_nsdi20, ; gavel_osdi20, ) over responsiveness and interactivity in that they allocate GPUs in a way that balances the competing demands of long-running jobs. Therefore, they are not well-suited for notebook IDLT. Whereas slight delays may have been acceptable before, even short delays can severely impact the user experience when performing highly interactive programming tasks. In NotebookOS, the primary goal is to provide users with fast GPU access for interactive sessions. This demands flexible and low-latency resource allocation.
@@
Design Choices. NotebookOS introduces several novel techniques and design decisions that harmonize between long-running, stateful notebook sessions and elastic GPU allocations for short-lived notebook tasks. These techniques include:
$\bullet$ Distributed Notebook Kernels: NotebookOS adopts a novel distributed notebook design, where each logical Jupyter kernel is mapped to a NotebookOS distributed kernel, each of which consists of $N$ kernel replicas scheduled across NotebookOS’s GPU server cluster. This distributed kernel is designed to oversubscribe the under-utilized GPU resources. Its benefit is twofold. (1) Replicated kernels increase the likelihood that at least one replica will have immediate access to GPU resources upon code submission (C2). (2) By oversubscribing GPUs, resource utilization can be improved significantly (C1).
@@
$\bullet$ Transparent GPU State Checkpointing: Instead of performing SMR, NotebookOS opts to checkpoint large GPU state objects (model parameters and training datasets) asynchronously and separately to a distributed storage system, anticipating that the destination replica may become active (C2).

### v2

# NotebookOS: A Replicated Notebook Platform for Interactive Training with On-Demand GPUs

@@ Abstract.
Interactive notebook programming is universal in modern ML and AI workflows, with interactive deep learning training (IDLT) emerging as a dominant use case. To ensure responsiveness, platforms like Jupyter and Colab reserve GPUs for long-running notebook sessions, despite their intermittent and sporadic GPU usage, leading to extremely low GPU utilization and prohibitively high costs. In this paper, we introduce NotebookOS, a GPU-efficient notebook platform tailored for the unique requirements of IDLT. NotebookOS employs replicated notebook kernels with Raft-synchronized replicas distributed across GPU servers. To optimize GPU utilization, NotebookOS oversubscribes server resources, leveraging high inter-arrival times in IDLT workloads, and allocates GPUs only during active cell execution. It also supports replica migration and automatic cluster scaling under high load. Altogether, this design enables interactive training with minimal delay. In evaluation on production workloads, NotebookOS saved over 1,187 GPU hours in 17.5 hours of real-world IDLT, while significantly improving interactivity.
Jupyter Notebook; Interactive Deep Learning Training; GPU Scheduling; Systems for AI
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3760250.3762230isbn: 979-8-4007-2165-6/2026/03ccs: Computer systems organization Cloud computingccs: Software and its engineering Schedulingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Traditionally, GPU-based, batch deep learning training (BDLT) features batch-style, long-running workloads (tiresias_nsdi19, ; gandiva_osdi18, ; optimus_eurosys18, ; themis_nsdi20, ; philly_trace_atc19, ; alibaba_cluster_2020_trace_nsdi22, ) that require uninterrupted access to GPU resources over long periods of time, such as large-scale model training. These traditional systems focus on maximizing throughput and job-completion time (JCT), often prioritizing long-running tasks that can afford latency and resource contention. In contrast, notebook-oriented, interactive deep learning training (IDLT) workloads—which consist of tasks like model and program debugging, dynamic/iterative model adjustments, hyperparameter tuning—demand a different set of performance characteristics. This set includes low latency, responsiveness, high interactivity, and efficient management of many short-running tasks. The short-lived tasks that dominate IDLT workloads are executed within the context of long-lived user sessions, such as Jupyter Notebook sessions.
@@
To guarantee interactivity, today’s notebook services typically provision and reserve GPU resources during the entire lifetime of a Notebook session. Although these sessions are long-running, cell executions are fragmented: IDLT workloads exhibit intermittent, sporadic, and often transient GPU usage. Users simply spend more time developing a notebook than they do executing the notebook’s tasks. As a result, notebook sessions spend a majority of their time not using the reserved GPUs, leading to extremely low GPU utilization and prohibitively high cost.
@@
We build NotebookOS, a first-of-its-kind notebook platform designed for the unique requirements of IDLT. NotebookOS is distinct in both purpose and design from traditional GPU cluster computing systems (gandiva_osdi18, ; themis_nsdi20, ; tiresias_nsdi19, ; gavel_osdi20, ; antman_osdi20, ; pollux_osdi21, ; sia_sosp23, ). NotebookOS prioritizes interactivity, low response time, and efficient management of numerous short-lived GPU tasks. By addressing these unique needs of users actively working in notebooks, NotebookOS realizes a novel notebook platform that uses fine-grained, flexible GPU allocation to effectively meet the demands of short, dynamic GPU tasks.
@@
At NotebookOS’s core is a collection of novel techniques and design tenets that enable the efficient support of IDLT workloads. NotebookOS uses a pluggable notebook task scheduling and placement mechanism, whose default policy is designed to maximize interactivity. Specifically, NotebookOS uses a replicated kernel design, where each notebook kernel is replicated across multiple GPU servers. Any replica can execute CPU or GPU tasks. Small kernel state is synchronized across replicas using Raft (raft_atc14, ), while large objects are asynchronously replicated via a distributed data store. This design decouples GPU allocation from notebook sessions: GPUs are dynamically allocated and assigned to one of the replicas only during code execution, maximizing the chances of immediate training upon code submission. To further improve GPU utilization, NotebookOS oversubscribes GPU resources on each server, leveraging the high task inter-arrival time (IAT) observed in IDLT workloads. This maximizes kernel availability and minimizes wait times during user code submission.
@@
    Comprehensive evaluation of NotebookOS using prototype and simulation on production IDLT workloads.
NotebookOS aims to pave the way for more efficient AI compute via on-demand GPUs. NotebookOS is available at: https://github.com/ds2-lab/NotebookOS.
@@ 2. Background and Motivation
This section presents an overview of relevant background and a workload analysis that motivates the need for a new platform tailored for notebook IDLT workloads.
Figure 1. Client-server Jupyter Notebook architecture.: Client-server Jupyter Notebook architecture.
@@ 2.1. Jupyter Notebook
Jupyter Notebook (jupyter_notebook, ) is an open-source, web-based interactive development environment (IDE) designed for interactive programming. Jupyter Notebook is most commonly used with Python and can be used with other languages (jupyter_available_kernels, ), such as R (jupyter_irkernel, ) and Julia (jupyter_ijulia, ). Jupyter Notebook has become the de facto development environment for data-driven programming: ML/AI, data science, and data analytics. This is due in large part to its versatility, ease-of-use, and interactivity.
Terminology. Notebooks are interactive documents that combine code, text for explaining the code, and visualizations for displaying the code’s inputs and outputs. These elements are contained within the cells of the notebook, which are the basic “units” of a notebook. When a user runs code in some cell(s), the code is sent to a separate process, called a kernel, which resides either on the user’s local computer or on a remote server. The kernel executes the code and returns results back to the notebook document where they are displayed in the cell(s). We refer to this process as a cell task execution. A notebook session is a persistent, working instance of a Jupyter Notebook environment where the state of variables, imports, and other execution context are maintained and reused by the associated kernel. Finally, we define an IDLT task as a cell task execution of GPU operations—such as model training and inference—during an interactive workload involving debugging and testing, initial model design, and other similar activities. See Figure 1 for an illustration.
@@ 2.2. Interactive Deep Learning Training
(a) Task duration CDF.
(b) Inter-arrival time CDF.
(c) GPU util. CDF (AdobeTrace).
(d) GPU & CPU usage (AdobeTrace).
Figure 2. Workload characteristics of three representative GPU cluster traces.: Workload characteristics of three representative GPU cluster traces.
@@ 2.3. Real-World IDLT Workload Analysis
Real-world IDLT workloads exhibit distinct behaviors compared to traditional long-running training jobs. In order to better understand these workloads and their requirements, we present a workload analysis of the production notebook IDLT traces collected from an Adobe internal research cluster (adobe, ). Adobe research cluster hosts containerized notebook sessions within AWS EC2 GPU virtual machine (VM) instances. Adobe research cluster also manages the GPU resource allocations for the notebook sessions. To deploy a notebook container, users specify how many GPUs should be allocated. Once launched, GPUs are bound to the notebook container, and the user can interact with the back-end notebook kernel (i.e., an IPython process) via a web-based Jupyter Notebook interface.
The IDLT traces (AdobeTrace) include 545,467 individual training events. We focus on a representative subset spanning June 1–August 31. The sample granularity of this trace is 15 seconds.
For context, we compare AdobeTrace with two other popular, publicly available GPU cluster traces. First, we use the Alibaba GPU Cluster’20 trace (alibaba_cluster_2020_trace_nsdi22, ) (AlibabaTrace), which captures both training and inference jobs running state-of-the-art ML algorithms from July to August 2020. Data were collected from a large production cluster with 6,500 GPUs across 1,800 servers. We also include the Philly GPU trace (philly_trace_atc19, ) (PhillyTrace), a 6.6GB representative subset of first-party deep learning training (DLT) workloads from Microsoft’s internal Philly clusters. PhillyTrace consists of 117,325 jobs collected from a cluster of 2,490 GPUs distributed across 552 servers between August 7 – December 22, 2017.
We chose these two traces for comparison because they represent typical DLT workloads, which are characterized by long-running training tasks that are scheduled by a batch scheduler and span several hours. As we show in our analysis, the workload captured by AdobeTrace differs considerably from that of PhillyTrace and AlibabaTrace.
@@ 2.3.1. Training Task Duration
[…]
The AdobeTrace workload consists of a significantly high fraction of short training tasks compared to both PhillyTrace and AlibabaTrace. Specifically, the 50th percentile of task durations is 120 seconds (2 minutes), 621 seconds (10.35 minutes), and 957 seconds (15.95 minutes) for AdobeTrace, PhillyTrace, and AlibabaTrace, respectively. This observation highlights the difference between traditional BDLT workloads (i.e., PhillyTrace and AlibabaTrace) and notebook IDLT workloads (i.e., AdobeTrace). Note that a meaningful comparison of task durations less than 15 seconds is impossible due to the 15-second granularity of AdobeTrace.
@@
2.3.2. Training Task Inter-Arrival Time (IAT)
Figure 2(b) shows a CDF plot of the IATs of training tasks from the Adobe research cluster trace. For all three traces, the IATs were measured within each user session independently, rather than the cluster-wide IATs of tasks submitted by any active user session, to ensure a fair comparison. From this graph, we can see that the 50th percentile of task IATs is 300 seconds (5 minutes), 44 seconds, and 38 seconds for AdobeTrace, PhillyTrace, and AlibabaTrace, respectively.
The observed AdobeTrace IATs align with the nature of IDLT workloads, where users intermittently make small, debugging-like changes to their models or Python code before submitting a short cell task to test the changes. In particular, AdobeTrace exhibits significantly longer IATs for more than 50% of tasks compared to PhillyTrace and AlibabaTrace. This is likely due to AdobeTrace users engaging in more extensive interactive cycles, iteratively testing and refining their code before each submission, rather than executing fully tested, production-ready DLT tasks scheduled by a batch GPU scheduler.
@@ 2.4. Insight and Challenges
This behavior leads to extremely low GPU utilization, as discussed in section˜2.3.
@@
$\bullet$ C3: Resource Elasticity. Low resource utilization stems from static GPU over-provisioning and a lack of elastic resource management in NaaS platforms. Adobe research cluster, operating roughly 12k V100 GPUs across 3k p3.16xlarge EC2 instances, incurs roughly \$18.3 million in monthly costs even with long-term reservation discounts. This high TCO (total cost of ownership), combined with the extremely low resource utilization, highlights the urgent need for a GPU-efficient notebook platform that can flexibly adapt based on real-time workload requirements.
@@ 3. NotebookOS Design
Workload Requirements. We present NotebookOS, a GPU-efficient notebook platform to address the challenges outlined in section˜2. Traditional GPU cluster schedulers are designed with long-running GPU-intensive tasks in mind. They often prioritize job completion time (JCT) (gandiva_osdi18, ; optimus_eurosys18, ; tiresias_nsdi19, ), throughput (gavel_osdi20, ), and fairness (themis_nsdi20, ; gavel_osdi20, ) over responsiveness and interactivity in that they allocate GPUs in a way that balances the competing demands of long-running jobs. Therefore, they are poorly-suited for notebook IDLT. Whereas slight delays may have been acceptable before, even short delays can severely impact the user experience when performing highly interactive tasks. In NotebookOS, the primary goal is to provide users with fast GPU access for interactive sessions. This demands flexible and low-latency resource allocation.
@@
Design Choices. NotebookOS introduces several novel techniques and design decisions that harmonize between long-running, stateful notebook sessions and elastic GPU allocations for short-lived notebook tasks, including:
$\bullet$ Distributed Notebook Kernels: NotebookOS adopts a novel distributed notebook design, where each logical Jupyter kernel is mapped to a NotebookOS distributed kernel, each of which consists of $N$ kernel replicas scheduled across NotebookOS’s GPU server cluster. This distributed kernel is designed to oversubscribe the under-utilized GPU resources. Its benefit is twofold. (1) Replicated kernels increase the likelihood that at least one replica will have immediate access to GPU resources upon code submission (C2). (2) Significantly improved resource utilization via GPU oversubscription (C1).
@@
$\bullet$ Transparent GPU State Checkpointing: Instead of performing SMR, NotebookOS checkpoints large GPU state objects (model parameters and training datasets) asynchronously and separately to a distributed storage system, anticipating that the destination replica may become active (C2).